# Network Modelling of Asynchronous Change-Points in Multivariate Time Series

### v1

@@ 2.2 Change-Point Prior
The change-point locations in each series are determined by the multivariate hidden state process $\{\boldsymbol{X}_{t}\}_{t=1,...,T}$ where $\boldsymbol{X}_{t}=(X_{1,t},...,X_{d,t})^{{}^{\prime}}$ is the vector containing the times of the most recent change-point before time $t$ for each series. Since the first segment in each series starts at time $t=1$, the process is initialised by
$X_{j,1}\overset{a.s.}{=}0\;\;\forall\;\;j=1,...,d$ (10)
and then for $t=2,...,T$ we have $\boldsymbol{X}_{t}\in\{0,...,t-1\}^{d}$, with the event $X_{j,t}=t-1$ indicating a change-point occurred at time $t$ in series $j$. Our prior over the hidden states assumes that the process $\{\boldsymbol{X}_{t}\}_{t=1,...,T}$ follows a Markov chain initialised by Equation (10) with time varying transition probabilities given by
@@
$\Pr(X_{j,t}=x_{j,t}|\boldsymbol{X}_{t-1}=\boldsymbol{x}_{t-1})=\begin{cases}p_{j,t}(\boldsymbol{x}_{t-1}),\;\;\;\;\;\;\;\;\;\;\;\;\text{if }x_{j,t}=t-1\\ 1-p_{j,t}(\boldsymbol{x}_{t-1}),\;\;\;\;\;\;\text{if }x_{j,t}=x_{j,t-1}\end{cases}$ (12)
@@
Conditional on $\boldsymbol{X}_{t-1}=\boldsymbol{x}_{t-1}$, the prior probability that a change-point occurs at time $t-1$ in series $j$ is given by $p_{j,t}(\boldsymbol{x}_{t-1})$. The dependence in the change-point locations across the series is controlled by our choice of $p_{j,t}(\boldsymbol{x}_{t-1})$.
@@
A natural starting point in specifying $p_{j,t}(\boldsymbol{x}_{t-1})$ is to consider independent univariate change-point priors for each series $j$. A popular choice is the Bernoulli process (see Yao,, 1984; Barry and Hartigan,, 1993; Fearnhead,, 2006), which specifies
$p_{j,t}(\boldsymbol{x}_{t-1})=q_{0,j}\;\;\;\text{for }j=1,...,d$ (13)
@@
for $q_{0,j}\in(0,1)$ which implies that the change-point probabilities in each series are constant for all $t$. This is equivalent to the change-point locations in each series $j$ being generated by a renewal process with geometrically distributed inter-arrival times where $q_{0,j}$ is the distribution parameter. Similar to the approach of Hallgren et al., (2024), we use the Bernoulli process as a starting point such that, in the case of no dependence, our prior defaults to Equation (13).
Motivated by lead-lag relationships in change-points in different series, if a given series $i$ leads series $j$, then $p_{j,t}(\boldsymbol{x}_{t-1})$ should depend on $x_{i,t-1}$. That is, the probability of a change-point occurring in series $j$ should depend on the distance to the most recent change-point in series $i$. As noted earlier, we represent this dependence via a weighted directed graph. Let $\boldsymbol{A}\in\{0,1\}^{d\times d}$ be an adjacency matrix, with elements $A_{i,j}$ controlling which pairs of series exhibit lead-lag relationships. Let $\boldsymbol{W}_{0}\in\mathbb{R}_{+}^{d}$ be a vector of positive weights, and $\boldsymbol{W}\in\mathbb{R}_{+}^{d\times d}$ a matrix of positive weights with elements $W_{i,j}$ controlling the strength of the lead-lag relationships. Finally, let $g_{i,j}:\mathbb{Z}_{+}\mapsto[0,1]$ be a set of pre-specified functions, with the function $g_{i,j}$ controlling the delay effect of a change-point in series $j$ following a change-point in series $i$.
Then our model for $p_{j,t}(\boldsymbol{x}_{t-1})$ with the latent weighted directed graph is given by:
$p_{j,t}(\boldsymbol{x}_{t-1})=Z_{j}^{-1}\left(W_{0,j}q_{0,j}+\sum_{i=1}^{d}A_{i,j}W_{i,j}g_{i,j}(t-x_{i,t-1}-1)I(x_{i,t-1}>0)\right)$ (14)
where $Z_{j}$ is a normalising constant given by
$Z_{j}=W_{0,j}+\sum_{i=1}^{d}A_{i,j}W_{i,j}$ (15)
and $I(x_{i,t-1}>0)$ the indicator function preventing the model from treating $t=0$ as an observed change-point. The normalised version of Equation (14) is given by:
$p_{j,t}(\boldsymbol{x}_{t-1})=W^{*}_{0,j}q_{0,j}+\sum_{\begin{subarray}{c}i=1\end{subarray}}^{d}W^{*}_{i,j}g_{i,j}(t-x_{i,t-1}-1)I(x_{i,t-1}>0)$ (16)
where $W^{*}_{0,j}=W_{0,j}/Z_{j}$ and $W^{*}_{i,j}=A_{i,j}W_{i,j}/Z_{j}$ with $\sum_{i=0}^{d}W_{i,j}^{*}=1$ are normalized weights. Equation (16) shows that the event of a change-point occurring at time $t$ in series $j$ is a mixture of $d+1$ Bernoulli trials. This mixture consists of the independent Bernoulli prior and a contribution from the lead-lag relationship between the $i$th series and $j$th series, (i.e. those that are connected to $j$ on the graph with $A_{i,j}=1$). The contribution of a connected series $i$ at a given time $t$ depends on the distance $(t-x_{i,t-1}-1)$ to the most recent change-point in that series.
From Equation (14) we can see that if $A_{i,j}=0\;\;\forall\;\;(i,j)$ or $d=1$ then we recover the independent Bernoulli prior (13) applied to each series. However, if $d>1$ and $A_{i,j}=1$ for some $(i,j)$, then the probability of a change-point occurring in series $j$ at a given time $t$ depends on the distance to the most recent change-point in series $i$. Since, $W_{i,j}$ and $g_{i,j}$ are strictly positive, a recent change in $i$ will cause an increase in the probability of a change occurring in $j$ shortly after. The magnitude of this increase is controlled by $W_{i,j}$ and the delay effect by $g_{i,j}$. Finally, the background weights $W_{0,j}$ control the contribution from the independent Bernoulli process (with parameter $q_{0,j}$) for each series $j$.
The choice of function $g_{i,j}$ depends on our prior beliefs regarding the delay in change-points between two dependent series. If we expect short delays, then we may choose a function which decays as the distance to the most recent change-point increases. A natural choice here is to use the probability mass function (PMF) of the geometric distribution:
@@
for some $q_{i,j}\in(0,1)$. Under this choice, change-points in a lagging series are most likely to occur shortly after those in the corresponding leading series. The delay can be controlled through the geometric decay parameters $\boldsymbol{q}=(q_{i,j})$, with a larger $q_{i,j}$ inducing a faster decay and shorter delays.
We assume $g_{i,j}$ is given by the geometric PMFs in Equation (17) and so our prior on the hidden states, $\boldsymbol{X}_{t}$, is parametrised by $(\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A},\boldsymbol{q}_{0},\boldsymbol{q})$. For brevity, we will suppress conditioning on $(\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A},\boldsymbol{q}_{0},\boldsymbol{q})$, implicitly assuming that any probabilities involving the hidden state process $\boldsymbol{X}_{t}$ depend on these parameters. In Section 2.3 we discuss how these parameters affect the properties of the hidden state process and hence the locations of the change-points.
@@ 2.3 Properties of the Prior
In this Section we assume that the hidden state hyperparameters $(\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A},\boldsymbol{q}_{0},\boldsymbol{q})$ are known and in Section 2.4 we discuss sensible hyperpriors for these. Recall that our prior on the hidden states is constructed such that when $\boldsymbol{A}=\boldsymbol{0}$, the change-points occur independently in each series according to a Bernoulli process with parameter $q_{0,j}$. When $\boldsymbol{A}\neq\boldsymbol{0}$, we have dependence between pairs of series $(i,j)$ for which $A_{i,j}\neq 0$.
A first step to understanding this dependence is to simulate a realisation of $\boldsymbol{X}_{t}$ and the corresponding probabilities $p_{j,t}$. Figure 1(a) displays a realisation of $\boldsymbol{X}_{t}$ for $d=3$. Here, $\boldsymbol{A}$ corresponds to a chain graph such that the second series depends on the first series and the third depends on the second. The change-points in the first series occur independently according to a Bernoulli process with parameter $q_{1,0}$. However, when a change occurs in the first series, the change probabilities in the second series experience an impulse response. This impulse response decays over time, due to the geometric function in Equation (17). This increases the likelihood that changes in the second series follow those in the first series shortly after, producing a lead-lag relationship in the change locations. The same effect is observed between series two and three, due to the directed edge between them.
1233
123
Figure 1: (a)–(b) Simulated change-point processes $\boldsymbol{X}_{t}$ with $d=3$, using a chain graph structure for $\boldsymbol{A}$ in (a) and a cyclic graph in (b). Bars represent the probabilities $p_{j,t}$, and dashed lines indicate the change-point locations. Parameters are the same for both cases: $\boldsymbol{q}_{0}=\frac{1}{50}\boldsymbol{1}_{3}$, $\boldsymbol{W}_{0}=\boldsymbol{1}_{3}$, $\boldsymbol{q}=\frac{1}{2}\boldsymbol{1}_{3\times 3}$, and $\boldsymbol{W}=\boldsymbol{1}_{3\times 3}$; (c)–(d) Cross-correlation function $Cor(U_{1,2},U_{j,2+h})$ for $j=1$ (top), $j=2$ (middle), and $j=3$ (bottom), with $\boldsymbol{A}$ as a chain graph in (c) and a cyclic graph in (d). Each column corresponds to a different edge weight $w$, and within each plot, curves show the effect of varying decay parameter $q$.
However, graphs that contain cycles can exhibit problematic dynamics, where bursts of consecutive change-points become likely due to excessive positive feedback. This effect is illustrated in Figure 1(b), where a change-point in the first series triggers an impulse response not only in the change-point probabilities of the second series two but also in its own. A subsequent change in the second series then feeds back into the first one, creating mutually reinforcing impulses. These feedback loops result in clusters of change-points occurring in quick succession. Such behaviour is typically undesirable because it leads to overfitting, where the model interprets random noise as meaningful structure, resulting in spurious change-point detections, reduced interpretability, and poor generalization to new data. Therefore, we should place little or no prior mass on cyclic graph structures, as further discussed in Section 2.4.
While the adjacency matrix $\boldsymbol{A}$ controls which pairs of series exhibit dependence, the degree of dependence is controlled by $(W_{i,j},q_{i,j})$. To quantify this dependence, we calculate the a priori correlation in the change-point locations at different lags for two given series. To do so we introduce a set of indicator variables $U_{j,t}$ defined as
$U_{j,t}=\begin{cases}1,\;\;\text{ if }X_{j,t}=t-1\\ 0,\;\;\text{ if }X_{j,t}=x_{j,t-1}\end{cases}$ (18)
i.e. $U_{j,t}=1$ indicates the presence of a change-point in series $j$ at time $t-1$. Then, for any two series $(i,j)$ we compute $\text{Cor}(U_{i,t},U_{j,t+h})$ for varying lags $h$. These correlations can be computed using a recursive calculation over the joint distribution of $\{\boldsymbol{X}_{s}\}_{s=1,...,t+h}$. In particular, for a given function $k(\cdot,\cdot)$, we can compute $\mathbb{E}(k(\boldsymbol{X}_{s},\boldsymbol{X}_{t}))$, for $t\geq s$, as
$\mathbb{E}(k(\boldsymbol{X}_{s},\boldsymbol{X}_{t}))$ $=\sum_{\boldsymbol{x}_{1}\in\mathcal{X}_{1}}\cdots\sum_{\boldsymbol{x}_{t}\in\mathcal{X}_{t}}Pr(\boldsymbol{X}_{1}=\boldsymbol{x}_{1},...,\boldsymbol{X}_{t}=\boldsymbol{x}_{t})\;k(\boldsymbol{x}_{s},\boldsymbol{x}_{t})$ (19)
$=\sum_{\boldsymbol{x}_{2}\in\mathcal{X}_{2}}Pr(\boldsymbol{X}_{2}=\boldsymbol{x}_{2}|\boldsymbol{X}_{1}=\boldsymbol{x}_{1})\;\cdots\sum_{\boldsymbol{x}_{t}\in\mathcal{X}_{t}}Pr(\boldsymbol{X}_{t}=\boldsymbol{x}_{t}|\boldsymbol{X}_{t-1}=\boldsymbol{x}_{t-1})\;k(\boldsymbol{x}_{s},\boldsymbol{x}_{t})$
where $\mathcal{X}_{t}=\{0,1,...,t-1\}^{d}$ is the support set of $\boldsymbol{X}_{t}$. The calculation in Equation (19) can easily be performed on a computer in low dimensions ($d<5$) and allows us to compute $\mathbb{E}(U_{i,t})$, $\mathbb{E}(U_{j,t+h})$, $\mathbb{E}(U_{i,t}U_{j,t+h})$ and hence $\text{Cor}(U_{i,t},U_{j,t+h})$ for varying $h$.
[…]
Figure 1(d) displays the correlations induced by the cyclic graph displayed in Figure 1(b). Focusing on the non-zero correlations in the first row, it is clear that the change-point indicators for the first series exhibit autocorrelation. This is expected and it reflects the positive feedback loops creating bursts of consecutive change-points which, as discussed previously, generate undesirable behaviour.

### v2

@@ 2.2 Change-Point Prior
We now describe our prior on the hidden state process, which we refer to as the NetCP process. The change-point locations in each series are determined by the multivariate hidden state process $\{\boldsymbol{X}_{t}\}_{t=1,...,T}$ where $\boldsymbol{X}_{t}=(X_{1,t},...,X_{d,t})^{{}^{\prime}}$ is the vector of run lengths for each series. Since the first segment in each series starts at time $t=1$, the process is initialized by
$X_{j,1}\overset{a.s.}{=}1\;\;\forall\;\;j=1,...,d$ (10)
and for t=2,,Tt=2,...,T we have $\boldsymbol{X}_{t}\in\{1,...,t\}^{d}$, with the event $X_{j,t}=1$ indicating a change-point occurred at time t1t-1 in series $j$. Our prior over the hidden states assumes that the process $\{\boldsymbol{X}_{t}\}_{t=1,...,T}$ follows a Markov chain initialized by Equation (10) with transition probabilities given by
@@
$\Pr(X_{j,t}=x_{j,t}|\boldsymbol{X}_{t-1}=\boldsymbol{x}_{t-1})=\begin{cases}p_{j,t}(\boldsymbol{x}_{t-1}),\;\;\;\;\;\;\;\;\;\;\;\;\text{if }x_{j,t}=1\\ 1-p_{j,t}(\boldsymbol{x}_{t-1}),\;\;\;\;\;\;\text{if }x_{j,t}=x_{j,t-1}+1\end{cases}$ (12)
@@
Conditional on $\boldsymbol{X}_{t-1}=\boldsymbol{x}_{t-1}$, the prior probability that a change-point occurs at time t1t-1 in series $j$ is given by $p_{j,t}(\boldsymbol{x}_{t-1})$. The dependence in the change-point locations across the series is controlled by our choice of $p_{j,t}(\boldsymbol{x}_{t-1})$.
@@
A natural starting point in specifying $p_{j,t}(\boldsymbol{x}_{t-1})$ is to consider independent univariate change-point priors for each series $j$. A popular choice is the Bernoulli process [[, see]]Yao1984, Barry1993, where
$p_{j,t}(\boldsymbol{x}_{t-1})=q_{0,j}\;\;\;\text{for }j=1,...,d\;\;\;\text{and }q_{0,j}\in(0,1),$ (13)
@@
implying that the change-point probabilities in each series are constant for all $t$. Here, the change-point locations in each series $j$ form a renewal process with geometrically distributed inter-arrival times where $q_{0,j}$ is the distribution parameter. Similar to the approach of [69], we use the Bernoulli process as a starting point such that, in the case of no dependence, our prior defaults to Equation (13).
Now, in order to incorporate lead-lag dependencies, if change-points in series $i$ lead those in series $j$, then $p_{j,t}(\boldsymbol{x}_{t-1})$ should depend on $x_{i,t-1}$. That is, the probability of a change-point occurring in series $j$ should depend on the distance (run length) to the most recent change-point in series $i$. We represent this dependence via a weighted directed graph, such that the presence of an edge $(i,j)$ indicates that change-points in $j$ are likely to follow change-points in $i$.
For each j=1,,dj=1,...,d, we refer to $q_{0,j}\in(0,1)$ in Equation (13), as the background rate of change-points in series $j$. We then let $W_{0,j}>0$ be the weight assigned to the background change-point process. To indicate which pairs of series $(i,j)$ exhibit lead-lag relationships, we define an adjacency matrix with entries $A_{i,j}\in\{0,1\}$. Then, for each edge $(i,j)$ we define an edge weight $W_{i,j}>0$ and an edge impulse function $g_{i,j}:\mathbb{Z}\rightarrow[0,1]$. Our model for $p_{j,t}(\boldsymbol{x}_{t-1})$ is then given by
$p_{j,t}(\boldsymbol{x}_{t-1})=Z_{j}^{-1}\left[W_{0,j}q_{0,j}+\sum_{i=1}^{d}A_{i,j}W_{i,j}g_{i,j}(x_{i,t-1})\boldsymbol{1}(x_{i,t-1}<t-1)\right]$ (14)
where $Z_{j}$ is a normalizing constant given by $Z_{j}=W_{0,j}+\sum_{i=1}^{d}A_{i,j}W_{i,j}$ and $\boldsymbol{1}(x_{i,t-1}<t-1)$ is the indicator function necessary for preventing the model treating $t=0$ as an observed change-point. We can also write the model in terms of a set of normalized weights as
$p_{j,t}(\boldsymbol{x}_{t-1})=W_{0,j}^{*}q_{0,j}^{*}+\sum_{i=1}^{d}W_{i,j}^{*}g(x_{i,t-1})\boldsymbol{1}(x_{i,t-1}<t-1)$ (15)
with $W_{0,j}^{*}=W_{0,j}/Z_{j}$ and $W_{i,j}^{*}=A_{i,j}W_{i,j}/Z_{j}$ for i=1,,di=1,...,d. Since we are interested in uncovering the latent network structure of the multivariate series we use the model representation in Equation (14). Doing so allows us to directly study the effect the edge adjacency matrix has on the lead-lag dynamics between the series.
Equation (14) presents the probability of series $j$ transitioning to a change-point, $p_{j,t}(\boldsymbol{x}_{t-1})$, as a weighted sum of the background rate $q_{0,j}$ and a linear combination positive impulses which depend on the distance to the most recent change-point in series $i$ such that $A_{i,j}=1$. The magnitude of each impulse is controlled by the edge weight $W_{i,j}$ and its shape controlled by the edge impulse function $g_{i,j}(x_{i,t-1})$. However, if $A_{i,j}=0$ for all i=1,,di=1,...,d then $p_{j,t}(\boldsymbol{x}_{t-1})$ defaults back to the independent Bernoulli process of Equation (13).
Due to the term $\boldsymbol{1}(x_{i,t-1}<t-1)$, introduced to prevent $t=0$ from being treated as a change point, our Markov process has inhomogeneous transitions. However, after all series have hit their first change-points, the chain is governed by the following homogeneous transitions
$p_{j}(\boldsymbol{x}_{t-1})=Z_{j}^{-1}\left[W_{0,j}q_{0,j}+\sum_{i=1}^{d}A_{i,j}W_{i,j}g_{i,j}(x_{i,t-1})\right]$ (16)
The choice of edge impulse function, $g_{i,j}(x)$, is an important one. This function is analogous to those seen in the network Hawkes process literature [77]. Functions with peaks close to one encourage change-points to occur shortly after those in leading series. Whereas functions with “humps” away from one can be used to model more distant dependence. In this article we restrict our attention to the case where we expect short delays between change-points. We therefore specify $g_{i,j}(t)$ to be the PMF of the geometric distribution
@@
where parameter $q_{i,j}\in(0,1)$ controls the rate of decay. Larger values of $q_{i,j}$ will result in a larger mass near one, promoting short delays, while values nearer to zero will flatten the function, encouraging longer delays.
The construction of $p_{j,t}(\boldsymbol{x}_{t-1})$ induces a form of mutual excitation between the change-point processes for each series. This is illustrated in Figure 1 which shows a realization of the prior for $d=2$ and $\boldsymbol{A}$ representing a chain graph with single edge $(1,2)$. Here, a change-point in series one causes an impulse in the change-point probabilities in series two. This impulse decays geometrically according to $g_{1,2}(\cdot)$ and encourages change-points to occur in series two shortly after.
122
Figure 1: Simulation from the NetCP process (right) for $d=2$ with $\boldsymbol{A}$ representing a chain graph (left) with $A_{1,2}=1$. Blue bars show the change-point probabilities, $p_{j,t}(\boldsymbol{x}_{t-1})$, and solid lines the change-point locations (where $X_{j,t}=1$). We use the parameters $q_{0,1}=q_{0,2}=0.1$, $W_{0,1}=W_{0,2}=1$, $W_{1,2}=W_{2,1}=5$ and $q_{1,2}=q_{2,1}=0.5$.
@@ 2.3 Properties of the Prior
In this section we examine key properties of our prior, including stationarity, the expected number of change-points and the implied cross correlation between the change-point locations in different series. A clear understanding of the properties of our prior model can guide our specification of hyperpriors for parameters $(\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A},\boldsymbol{q}_{0},\boldsymbol{q})$ that affect the lead-lag dependence of change points between series.
As noted in Section 2.2, the hidden state vector, $\boldsymbol{X}_{t}=(X_{1,t},...,X_{d,t})^{{}^{\prime}}$, is initially governed by inhomogeneous transitions given by Equation (14). These transitions imply
$W_{0,j}q_{0,j}/Z_{j}\leq p_{j,t}(\boldsymbol{x}_{t-1})<1,$
and so the rate at which the change points occur in series $j$ is greater than or equal to that of a Bernoulli process with parameter $W_{0,j}q_{0,j}/Z_{j}$. A trivial upper bound on the expected arrival time of the first change-point in series $j$ can then be obtained as
$\mathbb{E}(\tau_{j,1})\leq Z_{j}/W_{0,j}q_{0,j}$ (18)
Once all series experience their first change-point, the transitions become homogeneous, see Equation (16), and so investigating the limiting behavior of NetCP is equivalent to studying the behavior of the homogeneous phase.
Proposition 1.
Let $\boldsymbol{X}_{t}$ be a Markov process defined by the homogeneous transitions given by Equation (16). Then, for any initial condition, $\boldsymbol{X}_{t}$ converges to a unique stationary distribution.
Proposition 1 tells us that the homogeneous phase of the process is stationary and therefore does not exhibit degenerate or divergent behavior. Therefore, the limiting distribution of the NetCP process is stable and is equal to the stationary distribution of the homogeneous phase. The asymptotic stationarity of the chain guarantees that the run lengths, $\boldsymbol{X}_{t}=(X_{1,t},...,X_{d,t})^{{}^{\prime}},$ do not degenerate at 1 (constant change-points) or diverge to infinity (never having change-points).
However, the choice of graph structure and graph parameters $(\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A},\boldsymbol{q}_{0},\boldsymbol{q})$ can cause the chain to produce many change-points in a given time frame. To study the properties of the change-point locations we introduce a set of binary random variables
$U_{j,t}=\boldsymbol{1}(X_{j,t}=1)\in\{0,1\}$
indicating the locations where change-points occur in each series. We proceed by studying the behavior of the NetCP process when the structure of the graph is acyclic and cyclic.
2.3.1 Acyclic Graphs
In this subsection, we study the behavior of the NetCP when the underlying graph is acyclic. In order to build intuition for how the process behaves in this scenario, we consider the simple case of a chain graph with $d=2$ and a single edge $(1,2)$. In this scenario, we have $A_{1,1}=A_{2,1}=0$, and so the change-point process in series one evolves according to a Bernoulli process, independently of series two, with $Pr(U_{1,t}=1)=q_{0,1}$ for all $t>1$. However, the presence of the edge $(1,2)$ means that the rate of change-points in series two is dependent on series one.
Proposition 2.
Suppose $\boldsymbol{X}_{t}$ is the two-dimensional process governed by the NetCP transitions given in (14). Suppose that $\boldsymbol{A}$ represents a chain graph with $A_{1,2}=1$ and an edge impulse function, $g_{1,2}(.),$ given by Equation (17). Then for $t>1$
$Pr(U_{2,t}=1)=\frac{W_{0,2}q_{0,2}+W_{1,2}\lambda(t-1,q_{0,1},q_{1,2})}{W_{0,2}+W_{1,2}}$ (19)
where
$\lambda(t,q_{0,1},q_{1,2})=q_{0,1}q_{1,2}\left[\frac{1-[(1-q_{0,1})(1-q_{1,2})]^{t-1}}{1-(1-q_{0,1})(1-q_{1,2})}\right]$ (20)
From Proposition 2, we can see that the marginal probability of a change-point in series two at time $t$ is a weighted combination of the background rate for series two and $\lambda(t-1,q_{0,1},q_{1,2})$, which quantifies the contribution of the edge impulse function acting on the most recent change-point in series one. Note that $\lim_{q_{1,2}\rightarrow 1}\lambda(t,q_{0,1},q_{1,2})=q_{0,1}$, suggesting that as we make the edge impulse function more peaked, $Pr(U_{2,t}=1)$ becomes a simple weighted average of the background rates for the two series. For a sample of size $T$, the expected number of change-points in series two is then:
$\sum_{t=2}^{T}Pr(U_{2,t}=1)=\frac{(T-1)W_{0,2}q_{0,2}+W_{1,2}\sum_{t=2}^{T}\lambda(t-1,q_{0,1},q_{1,2})}{W_{0,2}+W_{1,2}}$ (21)
which remains a linear combination of the background rate for series two and the contribution from the edge impulse function from series one. The limiting (stationary) probability of a change-point in series two is then given by
$\lim_{t\rightarrow\infty}Pr(U_{2,t}=1)=\frac{W_{0,2}q_{0,2}+W_{1,2}\lambda(q_{0,1},q_{1,2})}{W_{0,2}+W_{1,2}}$ (22)
where $\lambda(q_{0,1},q_{1,2})=q_{0,1}q_{1,2}/[1-(1-q_{0,1})(1-q_{1,2})]$.
The key motivation for the construction of the NetCP process was to induce lead-lag dependence across the change-point locations in different series. To quantify this dependence we compute the cross-autocovariance and cross-autocorrelation in the change-point locations. This is formalized in the next proposition for the chain graph in $d=2$.
Proposition 3.
Suppose $\boldsymbol{X}_{t}$ is the two-dimensional process governed by the NetCP transitions given in Equation (14). Suppose that $\boldsymbol{A}$ represents a chain graph with $A_{1,2}=1$ and edge impulse function, $g_{1,2}(.),$ given by Equation (17). Then for $t>1$ and $h\geq 1$
$\text{Cov}(U_{1,t},U_{2,t+h})=\frac{q_{0,1}W_{1,2}[\lambda^{*}(h,q_{0,1},q_{1,2})-\lambda(t+h-1,q_{0,1},q_{1,2})]}{W_{0,2}+W_{1,2}}$ (23)
where $\lambda^{*}(t,q_{0,1},q_{1,2})=\lambda(t,q_{0,1},q_{1,2})+q_{12}[(1-q_{0,1})(1-q_{1,2})]^{t-1}$ and $\lambda(t,q_{0,1},q_{1,2})$ given in Proposition 2.
Combining Proposition 3 with $Pr(U_{2,t}=1)$ from Proposition 2 and $P(U_{1,t}=1)=q_{0,1}$ we obtain the cross-autocorrelation. The stationary cross-autocorrelation can then be calculated by replacing $\lambda(t+h-1,q_{0,1},q_{1,2})$ with its limit defined by $\lambda(q_{0,1},q_{1,2})$.
Figure 2: Stationary cross-autocorrelation of change-point indicators $\text{Cor}(U_{1,t},U_{2,t+h})$ under the NetCP process for $d=2$ with $A_{1,2}=1$ for $h=1$ (left), $h=2$ (middle) and $h=3$ (right). Each plot displays a heat-map of the correlation induced by varying edge weight $W_{1,2}$ and geometric edge impulse function parameter $q_{1,2}$. We set the parameters $q_{0,1}=q_{0,2}=0.1$ and $W_{0,2}=1$.
Figure 2 displays the stationary cross-autocorrelation, $\text{Cor}(U_{1,t},U_{2,t+h})$, of the NetCP process using a chain graph in $d=2$, for varying $h$ and $(q_{1,2},W_{1,2})$. Here, we can see that the cross-autocorrelation is highest at lag $h=1$ and decreases as the lag increases. At $h=1$, the correlation increases both in $q_{1,2}$ and in $W_{1,2}$. However, at lags $h=2$ and $h=3$, the highest correlation is achieved by setting lower values of $q_{1,2}$. Recall that $q_{1,2}$ controls the flatness of the geometric edge impulse function $g_{1,2}(x_{1,t-1})$. Smaller values of $q_{1,2}$ result in a flatter edge impulse function, which places more mass on change-points occurring at longer lags. Therefore when the edge impulse function is the PMF or a geometric distribution, smaller values of $q_{i,j}$ should be used to obtain higher correlation at longer lags between series $i$ and $j$.
2.3.2 Cyclic Graphs
The mutual excitation observed for acyclic graphs induces cross-autocorrelation between pairs of series, resulting in the lead-lag behavior we wish to model. Cyclic graphs, however, result in positive feedback loops, whereby a change-point in a given series $j$ can trigger a change in a dependent series $i$ which can in turn trigger another change in $j$. This effect is illustrated for $d=2$ in Figure 3 where the graph contains the cycle $\{(1,2),(2,1)\}$. Here, bursts of change-points become likely, whereby a change-point in series one causes an impulse in series two, which then feeds back into series one and vice versa.
122
Figure 3: Simulation of the NetCP process (right) for $d=2$ with $\boldsymbol{A}$ representing a cyclic graph (left) with $A_{1,2}=A_{2,1}=1$. Blue bars show the change-point probabilities, $p_{j,t}(\boldsymbol{x}_{t-1})$, and solid lines the change-point locations (where $X_{j,t}=1$). We use the parameters $q_{0,1}=q_{0,2}=0.1$, $W_{0,1}=W_{0,2}=1$, $W_{1,2}=W_{2,1}=5$ and $q_{1,2}=q_{2,1}=0.5$.
It is not hard to verify that if $\boldsymbol{A}_{\text{D}}$ is the adjacency matrix of the acyclic graph with single edge $(1,2)$ and $\boldsymbol{A}_{\text{C}}$ the cyclic graph with edges $\{(1,2),(2,1)\}$, then $\text{Cov}(U_{1,t},U_{1,t+h}|\boldsymbol{A}_{\text{D}})=0$ and $\text{Cov}(U_{1,t},U_{1,t+h}|\boldsymbol{A}_{\text{C}})>0$. So under the cyclic graph, the indicators within a given series are positively correlated, suggesting bursts of consecutive change-points are likely. Bursts of change-points, as seen in Figure 3, are generally undesirable since in many applications a structural change tends to persist for some time. A prior which encourages many change-points in quick succession may therefore be prone to overfitting. For this reason, we recommend avoiding cyclic graph structures as discussed in Section 2.4.